. A group of two persons is to be selected from two boys and two girls. What is the probability that the group will have

(a) two boys (b) one boy (c) only girls

Grade:

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
a)probability of group containing 2 boys=(2c2)/(4c2)=1/6.
b)probability of group containing 1 boy=(2c1*2c1)/(4c2)=2/3.
c)probability of group containing only girls=(2c2)/(4c2)=1/6.
Thanks and Regards
Shaik Aasif
askIITians faculty